Shift report 20011121 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 21 2001 - 07:58:51 EST

  • Next message: jnorris: "Paper Comments"

    ------------------------------------------------------------
    Date: Wed Nov 21 07:58:50 2001, Shift: 00:00-08:00
    Supervisor: james norris, with: 
    Log book#11 pages 19 - 20
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 00:00 previous beam stored & cogged 
          (at ~22:00 20-Nov-01)
       At 02:14 voted to dump ZDC=281<(1/3)*1200
       At 02:40 gap cleaning in blue ring
       At 02:53 MCR called to warn of beam dump
       At 02:57 quench link interlock
       At 03:54 ramping for hysteresis
       At 04:13 preparing for injection
       At 04:32 filling, preparing to ramp
       At 05:20 quench link interlock
       At 06:39 ramping for hysteresis
       At 07:02 preparing for injection
       At 07:48 quench link interlock
       
    
                                 trigger 6
    2. Run 5915 - 22:35-00:20       200K
       Run 5916 - 00:24-00:53     107.2K
       Run 5917 - 02:39-02:53       7.4K
    
    3. Noticed in brntuple plots for runs
       5910 - signal suppression in low channels
              for h5, h6, h13, h14, h16 of 
              ZDC page
       5914 - single count blips in high channels
              for all plots of
              TPC.NumSeq_Trigger_1_Only page
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 21 2001 - 07:59:37 EST